OpenStudy (anonymous):

find the area of the trapezoid with height of 8m and base 14m and 19m

OpenStudy (lgbasallote):

\[\large Area = \frac{h(b_1 + b_2)}{2}\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (lgbasallote):

h = height \(b_1\) and \(b_2\) are the bases

OpenStudy (anonymous):

.5 (14+19) = .5(33) =* 16.5 m*

OpenStudy (lgbasallote):

careful..don;t forget to multiply by the height ;)

OpenStudy (anonymous):

yes but I don't know where i keep messing up at

OpenStudy (lgbasallote):

substitute the values \[Area = \frac{8(14 + 19)}{2}\] you can divide 8 by 2 right there \[Area = 4(14 + 19)\]\[Area = 4(33)\]

OpenStudy (anonymous):

16.5 m * 8 = 132m

OpenStudy (lgbasallote):

now just multiply 4 x 33 and you have your answer ^_^

OpenStudy (anonymous):

ok I was multiplying the wrong number which was 19 instead of addding thanks